OpenStudy (anonymous):

Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

OpenStudy (anonymous):

i think its 64

OpenStudy (anonymous):

# of students = X # of adults = Y X + Y = 150 45Y + (28.50)X = 4770 Solve for X and Y.

OpenStudy (anonymous):

X + Y= 150 45X + 28.5Y= 4770 Now using systems of elimination, multiply the top equation by 45 (since it is a whole number), so you can eliminate one variable and solve for the other. 45X + 45Y = 6750 -(45X + 28.5Y= 4770) 16.5Y= 1980 Y= 120 If Y= 120, and X + Y= 150, then X= 30. There were 30 adults and 120 students. Simple enough. :)